OpenStudy (anonymous):

Show that the function F(x)=(x-a)^2 * (x-b)^2 + x takes on some value (a+b)/2 for some value x please helpp

OpenStudy (kinggeorge):

So we want to show that \[(x-a)^2 \cdot (x-b)^2 + x={a+b \over 2}\]for some value of x.

OpenStudy (anonymous):

im will guess mvt for this one am i right?

OpenStudy (kinggeorge):

That's what I was thinking as well.

OpenStudy (anonymous):

oh maybe not. hmm we have \[F(a)=a\] \[F(b)=b\]

OpenStudy (anonymous):

no its under intermediate value theoremm

OpenStudy (anonymous):

so maybe it is intermediate value theorem, since if \(a<b\) we have \[a< \frac{a+b}{2}<b\]

OpenStudy (anonymous):

ok so done F is continuous, must take on all values between \(F(a)\) and \(F(b)\) so must take on all values between \(a\) and \(b\)

OpenStudy (kinggeorge):

Looks right. And if \(b<a\) \[b<{a+b\over2}<a\]

OpenStudy (anonymous):

perfectt thanks alot guyss